OpenStudy (anonymous):

How would you find the determinant of : ⎡3 2⎤ ⎣6 4⎦ Without using a calculator?

OpenStudy (therealmeeeee):

ok You actually do it the same way for any size matrix: you pick a row and column, and multiply the value for each element in the row by the "cofactor" or "minor" (the smaller determinant found by eliminating the row and the column of the element that you're working with). | 5 2 0 0 -2 | | 0 1 4 3 2 | | 0 0 2 6 3 | | 0 0 3 4 1 | | 0 0 0 0 2 | This particular matrix is pretty easy, because it has a lot of zeros. So if you start with the last row, you get: 2 x | 5 2 0 0 | | 0 1 4 3 | | 0 0 2 6 | | 0 0 3 4 | Then, using the first column, you'd get: 2 x 5 x | 1 4 3 | | 0 2 6 | | 0 3 4 | Then again using the first column: 2 x 5 x 1 x | 2 6 | | 3 4 | 2 x 5 x 1 x (2 x 4 - 3 x 6) = -100

OpenStudy (therealmeeeee):

do this help

OpenStudy (anonymous):

@DestisaurusRex lol do not listen to this guy ^^^ he copies answers! he got it from here http://answers.yahoo.com/question/index?qid=20070123154335AAIVKZd

OpenStudy (therealmeeeee):

omg no idid not u just a stalker

OpenStudy (anonymous):

Yes you did.You have the same words plus you did the same thing to the other questions.

OpenStudy (anonymous):

And its obvious its not from you cause you dont even know basic grammar

OpenStudy (therealmeeeee):

omg girl no i didnt not stalker

OpenStudy (alekos):

she's right i'm afraid

OpenStudy (alekos):

leave maths to the real mathematicians

OpenStudy (alekos):

for a 2x2 determinant \[\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]\] it's ad - bc

OpenStudy (alekos):

no calculator required. should be able to do it in your head